Find the maximum and minimum values of the function g(0) interval [o. 7 2θ-4 sin(θ) on the Preview Minimum value-pi/3+2pi Maximum value O Preview
Given the function f(z) = 2e - List the x-coordinates of the critical values (enter DNE if none) DNE List the x-coordinates of the inflection points (enter DNE if none) DNE List the intervals over which the function is increasing or decreasing (use DNE for any empty intervals) Increasing on DNE Preview Decreasing on -1/5 *Preview List the intervals over which the function is concave up or concave down (use DNE for any empty intervals) Concave up on -1/5 Concave down on DNE *Preview Preview
Consider the function r()24r5160z derivative to determine cach of the following 15. Differentiate r and use the All intervals on which r is increasing. If there are more than one intervals, separate them by a comma. Use open intervals and exact values. r increases on: [0,1 Preview 5] All intervals on which ris decreasing, If there are more than one intervals, separate them by a comma. Use open intervals and exact values. r decreases on:「0,15] *Preview The value(s) of x at which r has a relative maximum. If there are more than one solutions, separate them by a comma. Use exact values r has relative maximum(s) at2 Preview The value(s) of x at which r has a relative minimum. If there are more than one solutions, separate them by a comma. Use exact values r has relative minimum(s) atx Preview
Given the function f(z) ""e22 Determine the open interval(s) where the function is concave up 0o,-1.707)Ur-0.29289 Preview Determine the open interval(s) where the function is concave down Preview (-1.70710678,-0.292891 Determine any points of inflection. (-0.29289,0.0477)
Answer #1

Given: g (θ) 29-4 sin (θ) Find the derivative of the given function g (θ) = 28-4sin(8-g(θ) = 2-4 cos (9) For critical values